OpenStudy (anonymous):

polar problem, r = 10 sin theta, change to cartesian equation, i multiplied both sides by r, r^2 = 10 r sin theta , x^2 + y^2 = 10 y

OpenStudy (anonymous):

oh i think i got a circle

OpenStudy (amistre64):

10 sin(t) is a rose maybe?

OpenStudy (anonymous):

subtract 10 y from both sides, complete the square

OpenStudy (anonymous):

no we started with r = 10 sin theta

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=polar {10sin%28t%29}

OpenStudy (anonymous):

multiply both sides by r

OpenStudy (amistre64):

its a circle offcentered

OpenStudy (anonymous):

r^2 = 10 r sin theta, so x^2 + y^2 = 10 y

OpenStudy (anonymous):

x^2 + y^2 - 10 y = 0, and complete the suqare

OpenStudy (amistre64):

that looks good

OpenStudy (anonymous):

how did you get a rose?

OpenStudy (amistre64):

i dint; i was trying to remember what it looked like and apparentlly forgot :)

OpenStudy (anonymous):

no problem :) thanks